the combustion of 1.00 mol of glucose, C6H12O6, releases 2820 kJ of heat. If 2.0g of glucose is burned in a calorimeter containing 1.0 kg of water, and the temperature increases by 3.5°C, what is the heat capacity of the calorimeter?
Answer #1

Q = m c ∆T
Q = quantity of heat in joules (J)
m = mass of the substance acting as the environment in
grams (g)
c = specific heat capacity (4.19 for H2O) in J/(g oC)
∆T = change in temperature = Tfinal - Tinitial in oC

2 gm glucose (Molar mass‎: ‎180.16) = 2 / 18.016 = 0.0111 Mole

0.0111 Mole can be released = 0.0111 x 2820 = 31.3 kJ

Energy gained by water

1000 x 4.19 x 3.5 = 14665 J or 14.665 Kj

The calorimeter got the rest:

31.3 - 14.665 = 16.635 kj

heat capacity of the calorimeter = 16.635 kj / 3.5 = 4.75 kj